# A concise tutorial review of Reverse Osmosis and Electrodialysis

### v5

@@ 6 Conclusions and Outlook
In this tutorial review, we presented theory for reverse osmosis (RO) and electrodialysis (ED), explaining how both technologies are based on the same fundamental transport theory. This is the solution-friction (SF) theory, and for ED we solved it in the absence of convection, thus we did not discuss pressures. We used SF theory for RO but then only described neutral solutes. Finally we solved SF theory for the osmosis experiment based on ions and a charged membrane, and we compared with experimental data. For ED we also developed new equations for Donnan equilibrium that extend the standard ideal expression. We present analytical equations for current efficiency, showing that this is a process parameter, not a membrane material property. For RO we summarized the literature for SF theory for neutral solutes including also the effect of concentration polarization. The general derivation we provided of SF theory also results in the twice-extended Nernst-Planck equation which is generally applicable in describing ion flow in reverse osmosis and nanofiltration of salt solutions.
Topics that we did not address in this tutorial review are first of all that both for RO and NF we must implement the Nernst-Planck equation for ions and a charged membrane in a full module calculation, and beyond that extend the theory from simple 1:1 salt solutions to multi-ionic solutions, also for electrodialysis. Even the addition of one extra type of anion or cation can significantly change the entire modeling framework. In addition, in real water sources also the protonation degree of ions must be considered, which depends on local pH. At high concentrations, ions also associate in ion pairs. These effects are relevant to study because for instance an ammonium ion is acted on by the electrical field, but the neutral ammonia species is not. Thus rejection of these ions is strongly pH-dependent. For very tiny pores, a related topic is the effective size of ions, that has an impact on their partitioning and their mobility within the membranes. Ions with a higher charge will be hydrated better, and are expected to be slower. State-of-the-art theory for simultaneous transport and reaction of ions (such as acid-base reactions between ions) assumes that these reactions are very fast, but it is interesting to investigate whether that is a correct assumption. Another important assumption is local electroneutrality in channels and in membranes. Especially in reverse osmosis with membranes as thin as 100 nm, it is important to know if possibly Poisson’s equation must be used to replace the assumption of local electroneutrality in the membrane.
@@
$f_{\mathrm{elec},j}$ free energy density of Coulombic interactions between ions J/m3
$f_{\text{F-m}}$ friction factor of water with the membrane matrix mol.s/m5
@@
[…]
@@
$\infty$ conditions outside of membrane, or salt concentration in solution for a $z:z$ salt
(*) The words ‘ion’ and ‘solute’ are interchangeably used in this table.
(**) These fluxes across the membrane are defined per unit geometrical ‘outer’ area of a membrane, i.e., they are defined as superficial velocities, not interstitial. This means they are not defined for instance per cross-sectional area of the pores only.

### v6

@@ 6 Conclusions and Outlook
In this tutorial review, we presented theory for reverse osmosis (RO) and electrodialysis (ED), explaining how both technologies are based on the same fundamental transport theory. This is the solution-friction (SF) theory, and for ED we solved it in the absence of convection, thus we did not discuss pressures. We used SF theory for RO but then only described neutral solutes. Finally we solved SF theory for the osmosis experiment based on ions and a charged membrane, and we compared with experimental data. For ED we also developed new equations for Donnan equilibrium that extend the standard ideal expression. We present analytical equations for current efficiency, showing that this is a process parameter, not a membrane material property. For RO we summarized literature for SF theory for neutral solutes including also the effect of concentration polarization. The general derivation we provided of SF theory also results in the twice-extended Nernst-Planck equation which is generally applicable in describing ion flow in reverse osmosis and nanofiltration of salt solutions.
Topics that we did not address in this tutorial review are first of all that both for RO and NF we must implement the Nernst-Planck equation for ions and a charged membrane in a full module calculation, and beyond that extend the theory from simple 1:1 salt solutions to multi-ionic solutions, as we also should for electrodialysis. Even the addition of one extra type of anion or cation can significantly change the entire modeling framework. In addition, in real water sources also the protonation degree of ions must be considered, which depends on local pH. At high concentrations, ions also associate in ion pairs. These effects are relevant to study because for instance an ammonium ion is acted on by the electrical field, but the neutral ammonia species is not. Thus rejection of these ions is strongly pH-dependent. For very tiny pores, a related topic is the effective size of ions, that has an impact on their partitioning and their mobility within the membranes. Ions with a higher charge will be hydrated better, and are expected to be slower. State-of-the-art theory for simultaneous transport and reaction of ions (such as acid-base reactions between ions) assumes that these reactions are very fast, but it is interesting to investigate whether that is a correct assumption. Another important assumption is local electroneutrality in channels and in membranes. Especially in reverse osmosis with membranes as thin as 100 nm, it is important to know if possibly Poisson’s equation must be used to replace the assumption of local electroneutrality in the membrane.
@@
$f_{\mathrm{cou},j}$ free energy density of Coulombic interactions between ions J/m3
$f_{\text{w-m}}$ friction factor of water with the membrane matrix mol.s/m5
@@
$k_{\text{w-m}}$ water-membrane permeability, $=1/\left(f_{\text{w-m}}L\mathrm{{}_{m}}\right)$ m4/mol.s
@@
$\infty$ conditions outside membrane; salt concentration in solution for $z\!:\!z$ salt
(*) The words ion and solute are interchangeably used in this table.
(**) These fluxes across the membrane are defined per unit geometrical, i.e., outer, area of a membrane, i.e., they are defined as superficial velocities, not interstitial. This means they are not defined for instance per cross-sectional area of the pores only.